# Frame Interpolation with Consecutive Brownian Bridge Diffusion

### v2

@@ Abstract.
Recent work in Video Frame Interpolation (VFI) tries to formulate VFI as a diffusion-based conditional image generation problem, synthesizing the intermediate frame given a random noise and neighboring frames. Due to the relatively high resolution of videos, Latent Diffusion Models (LDMs) are employed as the conditional generation model, where the autoencoder compresses images into latent representations for diffusion and then reconstructs images from these latent representations. Such a formulation poses a crucial challenge: VFI expects that the output is deterministically equal to the ground truth intermediate frame, but LDMs randomly generate a diverse set of different images when the model runs multiple times. The reason for the diverse generation is that the cumulative variance (variance accumulated at each step of generation) of generated latent representations in LDMs is large. This makes the sampling trajectory random, resulting in diverse rather than deterministic generations. To address this problem, we propose our unique solution: Frame Interpolation with Consecutive Brownian Bridge Diffusion. Specifically, we propose consecutive Brownian Bridge diffusion that takes a deterministic initial value as input, resulting in a much smaller cumulative variance of generated latent representations. Our experiments suggest that our method can improve together with the improvement of the autoencoder and achieve state-of-the-art performance in VFI, leaving strong potential for further enhancement.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer vision
Figure 1. Qualitative Comparison of our proposed method and recent state-of-the-art methods. Overlaid Inputs are the average of two neighboring frames, and the results predict their intermediate frame. Our method generates better and clearer interpolation results than recent SOTAs, such as clearer dog skins (first row), clearer cloth with folds (second row), and clearer fences with nets and high-quality shoes (third row). Images within blue boxes are displayed to better compare detailed qualities, and red circles highlight our better performance. Examples are chosen from SNU-FILM (Choi et al., 2020) extreme subset which is the hardest one with large motion changes. More visual results are provided in Appendix C.2.
@@ 1. Introduction
Current works in VFI can be divided into two folds in terms of methodologies: flow-based methods (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021b; Dutta et al., 2022; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023b) and kernel-based methods (Chen et al., 2021; Cheng and Chen, 2020; Lee et al., 2020; Niklaus et al., 2017a, b; Shi et al., 2021; Dai et al., 2017). Flow-based methods compute flows in the neighboring frames and forward warp neighboring images and features (Hu et al., 2022; Niklaus and Liu, 2018, 2020; Jin et al., 2023; Siyao et al., 2021) or estimate flows from the intermediate frame to neighboring frames and backward warp neighboring frames and features (Lu et al., 2022a; Plack et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Dutta et al., 2022; Choi et al., 2021b; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023b). Instead of relying on optical flows, kernel-based methods predict convolution kernels for pixels in the neighboring frames. Recent advances in flow estimation (Huang et al., 2022a; Weinzaepfel et al., 2023; Hui et al., 2018; Huang et al., 2022c; Ilg et al., 2017; Teed and Deng, 2020; Sun et al., 2018) make it more popular to adopt flow-based methods in VFI.
Other than these two folds of methods, MCVD (Voleti et al., 2022) and LDMVFI (Danier et al., 2024) start formulating VFI as a diffusion-based image generation problem. LDMVFI considers VFI as a conditional generation task with Latent Diffusion Models (LDMs) (Rombach et al., 2022), where LDMs contain an autoencoder that compresses images into latent representations and reconstructs images from latent representations. Diffusion models (Ho et al., 2020) run in the latent space of the autoencoder. Though diffusion models achieve excellent performance in image generation, there remain challenges in applying them to VFI.
@@
    The formulation of diffusion models results in a large cumulative variance (the variance accumulated during sampling) of generated latent representations. The sampling process starts with standard Gaussian noise and adds small Gaussian noise to the denoised output at each step based on a pre-defined distribution. After the sampling process, images are generated, but these noises also add up to a large cumulative variance. Though such a variance is beneficial to diversity (i.e. repeated sampling results in different outputs), VFI requires that repeated sampling returns identical results, which is the ground truth intermediate frame. Therefore, a small cumulative variance is preferred in VFI. The relation of the cumulative variance and diversity is supported by the fact that DDIM (Song et al., 2021a) tends to generate relatively deterministic images than DDPM (Ho et al., 2020). DDIM removes small noises at each sampling step, so the cumulative variance in DDIM is lower. LDMVFI (Danier et al., 2024) uses conditional generation as guidance, but this does not change the nature of large cumulative variance. In Section 3.4, we show that our method has a much lower cumulative variance than conditional generation.
@@
    Videos usually have high resolution, which can be up to 4K (Perazzi et al., 2016), resulting in practical constraints to apply diffusion models (Ho et al., 2020) in pixel spaces. It is natural to apply Latent Diffusion Models (LDMs) (Rombach et al., 2022) to sample latent representations and reconstruct them back to images. LDMs apply VQModels in VQGAN (Esser et al., 2021) to compress images into latent representations and reconstruct images from latent representations. However, it does not take advantage of neighboring frames, which can be a good guide to reconstruction. LDMVFI designs reconstruction models that leverage neighboring frames, but it tends to reconstruct overlaid images when there is a relatively large motion between neighboring frames, possibly due to the cross-attention with features of neighboring frames, which is shown in Figure 1.
To tackle these challenges, we propose a consecutive Brownian Bridge diffusion model (in latent space) that transits among three deterministic endpoints for VFI. This method results in a much smaller cumulative variance, achieving a better estimation of the ground truth inputs. We can separate LDM-based VFI methods into two parts: autoencoder and ground truth estimation (with diffusion). It is different from the original LDMs (Rombach et al., 2022) because the latent representation generated by diffusion does not aim to estimate some ground truth. It is also different from LDMVFI (Danier et al., 2024) because LDMVFI does not consider the performance of autoencoder separately from the interpolation method. With such a two-stage separation, we evaluate them separately for specific directions of improvement. Moreover, we take advantage of flow estimation and refinement methods in recent literature (Lu et al., 2022a) to improve the autoencoder. The feature pyramids from neighboring frames are warped based on estimated optical flows, aiming to alleviate the issues of reconstructing overlaid images. In experiments, our method improves by a large margin when the autoencoder is improved and achieves state-of-the-art performance. Our contribution can be summarized in three parts:
  • We propose a new consecutive Brownian Bridge diffusion model for VFI and justify its advantages over traditional diffusion models: lower cumulative variance and better ground truth estimation capability. Additionally, we provide a cleaner formulation of Brownian Bridges and also propose the loss weights among different times in Brownian Bridges.
  • We formulate the diffusion-based VFI as two stages: autoencoder and ground truth estimation. This is a novel interpretation of LDM-based VFI, which can provide specific directions for improvements.
@@ 2.1. Video Frame Interpolation
Video Frame Interpolation can be roughly divided into two categories in terms of methodologies: flow-based methods (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021b; Dutta et al., 2022; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023b) and kernel-based methods (Chen et al., 2021; Cheng and Chen, 2020; Lee et al., 2020; Niklaus et al., 2017a, b; Shi et al., 2021; Dai et al., 2017). Flow-based methods assume certain motion types, where a few works assume non-linear types (Choi et al., 2021b; Dutta et al., 2022) while others assume linear. Via such assumptions, flow-based methods estimate flows in two ways. Some estimate flows from the intermediate frame to neighboring frames (or the reverse way) and apply backward warping to neighboring frames and their features (Lu et al., 2022a; Plack et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Dutta et al., 2022; Choi et al., 2021b; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023b). Others compute flows among the neighboring frames and apply forward splatting (Hu et al., 2022; Niklaus and Liu, 2018, 2020; Jin et al., 2023; Siyao et al., 2021). In addition to the basic framework, advanced details such as recurrence of inputs with different resolution level (Jin et al., 2023), cross-frame attention (Zhang et al., 2023b), and 4D-correlations (Li et al., 2023b) are proposed to improve performance. Kernel-based methods, introduced by (Niklaus et al., 2017a), aim to predict the convolution kernel applied to neighboring frames to generate the intermediate frame, but it has difficulty in dealing with large displacement. Following works (Dai et al., 2017; Cheng and Chen, 2020; Lee et al., 2020) alleviate such issues by introducing deformable convolution. LDMVFI (Danier et al., 2024) recently introduced a method based on Latent Diffusion Models (LDMs) (Rombach et al., 2022), formulating VFI as a conditional generation task. LDMVFI uses an autoencoder introduced by LDMs to compress images into latent representations, efficiently run the diffusion process, and then reconstruct images from latent space. Instead of directly predicting image pixels during reconstruction, it takes upsampled latent representations in the autoencoder as inputs to predict convolution kernels in kernel-based methods to complete the VFI task.
@@ 2.2. Diffusion Models
The diffusion model is introduced by DDPM (Ho et al., 2020) to image generation task and achieves excellent performance in high-fidelity and high-diversity image generation. The whole diffusion model can be split into a forward diffusion process and a backward sampling process. The forward diffusion process is defined as a Markov Chain with steps $t=1,...,T$, and the backward sampling process aims to estimate the distribution of the reversed Markov chain. The variance of the reversed Markov chain has a closed-form solution, and and expected value of the reversed Markov chain is estimated with a deep neural network. Though achieving strong performance in image generation tasks, DDPM (Ho et al., 2020) requires $T=1000$ iterative steps to generate images, resulting in inefficient generation. Sampling steps cannot be skipped without largely degrading performance because the conditional distribution at step $t-2$ needs to be computed with the conditional distribution at time $t-1$ and $t$ due to its Markov property. To enable efficient and high-quality generation, DDIM (Song et al., 2021a) proposes a non-Markov formulation of diffusion models, where the conditional distribution at time $t-k$ ($k>0$) can be directly computed with the conditional distribution at time $t$. Therefore, skipping steps does not largely degrade performance. Score-based SDEs (Batzolis et al., 2021; Song et al., 2021b; Zhou et al., 2024) are also proposed as an alternative formulation of diffusion models by writing the diffusion process in terms of Stochastic Differential Equations (Oksendal, 2013), where the reversed process has a closed-form continuous time formulation and can be solved with Eluer’s method with a few steps (Song et al., 2021b). In addition, Probability Flow ODE is proposed as the deterministic process that shares the same marginal distribution with the reversed SDE (Song et al., 2021b). Following score-based SDEs, some works propose efficient methods to estimate the solution Probability Flow ODE (Lu et al., 2022b, c). Instead of focusing on the nature of the diffusion process, DeepCache (Ma et al., 2024) proposes a feature caching and sharing mechanism in the denoising UNet, enabling parallel and skipping computation and further improving efficiency. To deal with high-resolution images, the Latent Diffusion Model (Rombach et al., 2022) proposes an autoencoder with a Vector Quantization Layer (VQ Layer) that compresses and reconstructs images, and diffusion models run with compressed images. With such an autoencoder, high-resolution images can be generated efficiently. Other than accelerating generation, diffusion models are applied to conditional generation tasks (Rombach et al., 2022; Zhang et al., 2023a; Batzolis et al., 2021; Choi et al., 2021a; Saharia et al., 2022; Zhou et al., 2024; Li et al., 2023a) such as generation based on poses or skeletons, image inpainting, etc.
@@ 3. Methodology
In this section, we will first go through preliminaries on the Diffusion Model (DDPM) (Ho et al., 2020) and Brownian Bridge Diffusion Model (BBDM) (Li et al., 2023a) and introduce the overview of our two-stage formulation: autoencoder and ground truth estimation (with consecutive Brownian Bridge diffusion). Then, we will discuss the details of our autoencoder method. Finally, we propose our solution to the frame interpolation task: consecutive Brownian Bridge diffusion.
@@ 3.1. Preliminaries
When $t=1$, $\mathbf{x}_{t-1}=\mathbf{x}_{0}$ is a sampled from the data (images). By iterating Eq. (1), we get the conditional marginal distribution of $\mathbf{x}_{t}$ (Ho et al., 2020):
@@
The sampling process can be derived with the Bayes’ theorem (Ho et al., 2020):
@@
Brownian Bridge Diffusion Model. Brownian Bridge (Ross, 1995) is a stochastic process that transits between two fixed endpoints, which is formulated as $X_{t}=W_{t}|(W_{t_{1}},W_{t_{2}})$, where $W_{t}$ is a standard Wiener process with distribution $\mathcal{N}(0,t)$. We can write a Brownian Bridge as $X_{t}=W_{t}|(W_{0},W_{T})$ to define a diffusion process. When $W_{0}=a,W_{T}=b$, it follows a normal distribution:
@@ 3.2. Formulation of Diffusion-based VFI
The goal of video frame interpolation is to estimate the intermediate frame $I_{n}$ given the previous frame $I_{0}$ and the next frame $I_{1}$. n is set to 0.5 to interpolate the frame in the middle of $I_{0}$ and $I_{1}$. In latent diffusion models (Rombach et al., 2022), there is an autoencoder that encodes images to latent representations and decodes images from latent representations. The diffusion model is given a standard Gaussian noise, denoises it according to the sampling process, and decodes the denoised latent representation back to an image. Since the initial noise is random, the decoded images are diverse images when they are sampled repetitively with the same conditions such as poses. Instead of diversity, VFI looks for a deterministic ground truth, which is the intermediate frame. Such a ground truth frame is encoded to a ground truth latent representation by the encoder, and only the ground truth latent representation needs to be estimated since the decoder will decode it back to the frame. Therefore, LDM-based VFI can be split into two stages: autoencoder and ground truth estimation. The two stages are defined as:
@@
The autoencoder is modeled with VQModel (Rombach et al., 2022) in Section 3.3, and the ground truth estimation is accomplished by our proposed (latent) consecutive Brownian Bridge diffusion in Section 3.4. During inference, both stages are combined as shown in Figure 2 (c), where we decode diffusion-generated latent representation $\hat{\mathbf{x}}$. Via such formulation, we can have a more specific direction to improve VFI quality. If images decoded from $\mathbf{x}$ (Figure 2 (a)) have similar visual quality to images decoded from $\hat{\mathbf{x}}$ (Figure 2 (c)), then the diffusion model achieves a strong performance in ground truth estimation, so it will be good to develop a good autoencoder. On the other way round, the performance of ground truth estimation can be potentially improved by redesigning the diffusion model.

### v3

@@ Abstract.
Recent work in Video Frame Interpolation (VFI) tries to formulate VFI as a diffusion-based conditional image generation problem, synthesizing the intermediate frame given a random noise and neighboring frames. Due to the relatively high resolution of videos, Latent Diffusion Models (LDMs) are employed to run diffusion models in latent space efficiently. Such a formulation poses a crucial challenge: VFI expects that the output is deterministically equal to the ground truth intermediate frame, but LDMs randomly generate a diverse set of different images when the model runs multiple times. The diversity is due to the large cumulative variance (variance accumulated at each generation step) of generated latent representations in LDMs, making the sampling trajectory random. To address this problem, we propose our unique solution: Frame Interpolation with Consecutive Brownian Bridge Diffusion. Specifically, we propose consecutive Brownian Bridge diffusion that takes a deterministic initial value as input, resulting in a much smaller cumulative variance of generated latent representations. Our experiments suggest that our method can improve together with the improvement of the autoencoder and achieve state-of-the-art performance in VFI, leaving strong potential for further enhancement. Our code is available at https://github.com/ZonglinL/ConsecutiveBrownianBridge.
@@
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.368096conference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiaisbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision
Figure 1. Qualitative Comparison of our proposed method and SOTAs. Our method generates clear interpolated images, while recent SOTAs generate blurred or overlaid results. Images within blue boxes are displayed for detailed comparisons, and red circles highlight our performance. Examples are chosen from SNU-FILM (Choi et al., 2020) extreme subset. More visualization results are shown in Appendix C.2
@@ 1. Introduction
Current works in VFI can be divided into two folds in terms of methodologies: flow-based methods (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021; Dutta et al., 2022; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023) and kernel-based methods (Chen et al., 2021; Cheng and Chen, 2020; Lee et al., 2020; Niklaus et al., 2017a, b; Shi et al., 2021; Dai et al., 2017). Flow-based methods take advantages of optical flow estimations to warp images and features to synthesize intermediate frames. Instead of relying on optical flows, kernel-based methods predict convolution kernels for pixels in the neighboring frames. Recent advances in flow estimation (Huang et al., 2022a; Weinzaepfel et al., 2023; Hui et al., 2018; Huang et al., 2022c; Ilg et al., 2017; Teed and Deng, 2020; Sun et al., 2018) make it more popular to adopt flow-based methods in VFI.
Other than these two folds of methods, MCVD (Voleti et al., 2022) and LDMVFI (Danier et al., 2024) start formulating VFI as a diffusion-based image generation problem, where LDMVFI takes advantage of LDM (Rombach et al., 2022) for better efficiency. Though diffusion models achieve excellent performance in image generation, there remain challenges in applying them to VFI.
@@
    The formulation of diffusion models results in a large cumulative variance (the variance accumulated during sampling) of generated latent representations. The sampling process starts with standard Gaussian noise and adds small Gaussian noise to the denoised output at each step. Noises are added up to a large cumulative variance when images are generated. Though such a variance is beneficial to diversity (i.e. repeated sampling results in different outputs), VFI requires that repeated sampling returns identical results, which is the ground truth intermediate frame. Therefore, a small cumulative variance is preferred in VFI. The relation of the cumulative variance and diversity is supported by the fact that DDIM (Song et al., 2021a) tends to generate relatively deterministic images than DDPM (Ho et al., 2020) because DDIM removes small noises at each sampling step. In Section 3.4, we show that our method has a much lower cumulative variance.
@@
    Videos usually have high resolution, which can be up to 4K (Perazzi et al., 2016), resulting in practical constraints to apply diffusion models (Ho et al., 2020) in pixel spaces. It is natural to apply Latent Diffusion Models (LDMs) (Rombach et al., 2022) for better efficiency, but this does not take advantage of neighboring frames, which can be a good guide to reconstruction. LDMVFI(Danier et al., 2024) designs reconstruction models that leverage neighboring frames, but it tends to reconstruct overlaid images when there is a relatively large motion between neighboring frames, possibly due to the cross-attention with features of neighboring frames, which is shown in Figure 1.
To tackle these challenges, we propose a consecutive Brownian Bridge Diffusion (in latent space) that transits among three deterministic endpoints for VFI. This method results in a much smaller cumulative variance, achieving a better estimation of the ground truth inputs. We also provide a novel method to analyze the LDM-based VFI methods: by analyzing the gap between quantitative metrics of the outputs from the autoencdoer and the outputs from the diffusion model + decoder (we name this as ground truth estimation), it is easier to figure out the specific directions of improvement. Moreover, we take advantage of flow estimation and refinement methods in recent literature (Lu et al., 2022a) to improve the autoencoder. The feature pyramids from neighboring frames are warped based on estimated optical flows, aiming to alleviate the issues of reconstructing overlaid images. In experiments, our method improves by a large margin when the autoencoder is improved and achieves state-of-the-art performance. Our contribution can be summarized in three parts:
  • We propose a novel consecutive Brownian Bridge diffusion model for VFI and justify its advantages over traditional diffusion models: lower cumulative variance and better ground truth estimation capability. Additionally, we provide a cleaner formulation of Brownian Bridges and also propose the loss weightng for Brownian Bridges.
  • We provide a novel method to analyze LDM-based VFI. With our methods of analysis, researchers can have specific directions for further improvements.
@@ 2.1. Video Frame Interpolation
Video Frame Interpolation can be roughly divided into two categories in terms of methodologies: flow-based methods (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021; Dutta et al., 2022; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023) and kernel-based methods (Chen et al., 2021; Cheng and Chen, 2020; Lee et al., 2020; Niklaus et al., 2017a, b; Shi et al., 2021; Dai et al., 2017). Flow-based methods assume certain motion types, where a few works assume non-linear types (Choi et al., 2021; Dutta et al., 2022) while others assume linear. Via such assumptions, flow-based methods estimate flows in two ways. Some estimate flows from the intermediate frame to neighboring frames (or the reverse way) and apply backward warping to neighboring frames and their features (Lu et al., 2022a; Plack et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Dutta et al., 2022; Choi et al., 2021; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023). Others compute flows among the neighboring frames and apply forward splatting (Hu et al., 2022; Niklaus and Liu, 2018, 2020; Jin et al., 2023; Siyao et al., 2021). In addition to the basic framework, advanced details such as recurrence of inputs with different resolution level (Jin et al., 2023), cross-frame attention (Zhang et al., 2023), and 4D-correlations (Li et al., 2023b) are proposed to improve performance. Kernel-based methods, introduced by (Niklaus et al., 2017a), aim to predict the convolution kernel applied to neighboring frames to generate the intermediate frame, but it has difficulty in dealing with large displacement. Following works (Dai et al., 2017; Cheng and Chen, 2020; Lee et al., 2020) alleviate such issues by introducing deformable convolution. LDMVFI (Danier et al., 2024) recently introduced a method based on Latent Diffusion Models (LDMs) (Rombach et al., 2022), formulating VFI as a conditional generation task. LDMVFI uses an autoencoder introduced by LDMs to compress images into latent representations, efficiently run the diffusion process, and then reconstruct images from latent space. Instead of directly predicting image pixels during reconstruction, it takes upsampled latent representations in the autoencoder as inputs to predict convolution kernels in kernel-based methods to complete the VFI task.
@@ 2.2. Diffusion Models
The diffusion model is introduced by DDPM (Ho et al., 2020) to image generation task and achieves excellent performance in image generation. The whole diffusion model can be split into a forward diffusion process and a backward sampling process. The forward diffusion process is defined as a Markov Chain with steps $t=1,...,T$, and the backward sampling process aims to estimate the distribution of the reversed Markov chain. The variance of the reversed Markov chain has a closed-form solution, and the expected value is estimated with a deep neural network. Though achieving strong performance in image generation tasks, DDPM (Ho et al., 2020) requires $T=1000$ iterative steps to generate images, resulting in inefficient generation. Sampling steps cannot be skipped without largely degrading performance due to its Markov chain property. To enable efficient and high-quality generation, DDIM (Song et al., 2021a) proposes a non-Markov formulation of diffusion models, where the conditional distribution at time $t-k$ ($k>0$) can be directly computed with the conditional distribution at time $t$. Therefore, skipping steps does not largely degrade performance. Score-based SDEs (Batzolis et al., 2021; Song et al., 2021b; Zhou et al., 2024) are also proposed as an alternative formulation of diffusion models by writing the diffusion process in terms of Stochastic Differential Equations (Oksendal, 2013), where the reversed process has a closed-form continuous time formulation and can be solved with Eluer’s method with a few steps (Song et al., 2021b). In addition, Probability Flow ODE is proposed as the deterministic process that shares the same marginal distribution with the reversed SDE (Song et al., 2021b). Following score-based SDEs, some works propose efficient methods to estimate the solution Probability Flow ODE (Lu et al., 2022b, c). Other than using the diffusion process to connect data distribution and Gaussian distribution, diffusion bridges (Li et al., 2023a; Zhou et al., 2024; De Bortoli et al., 2021; Shi et al., 2024) are proposed to connect arbitrary distributions such as two different data distributions. Instead of working on the diffusion process, the Latent Diffusion Model (Rombach et al., 2022) proposes autoencoders with KL-regularized (VAE) and VQ-regularized (VQ Layer) that compress and reconstruct images, and diffusion models run with compressed images. With such autoencoders, high-resolution images can be generated efficiently. In our work, the Vector Quantized (VQ) version is deployed.
@@ 3. Methodology
In this section, we will first go through preliminaries on the Diffusion Model (DDPM) (Ho et al., 2020) and Brownian Bridge Diffusion Model (BBDM) (Li et al., 2023a) and introduce the overview of the two-stage formulation: autoencoder and ground truth estimation (with consecutive Brownian Bridge diffusion). Then, we will discuss the details of our autoencoder method. Finally, we propose our solution to the frame interpolation task: consecutive Brownian Bridge diffusion.
@@ 3.1. Preliminaries
When $t=1$, $\mathbf{x}_{t-1}=\mathbf{x}_{0}$ is a sampled from the data (images). By iterating Eq. (1), we get the conditional marginal distribution (Ho et al., 2020):
@@
The sampling process is derived with the Bayes’ theorem (Ho et al., 2020):
@@
Brownian Bridge Diffusion Model. Brownian Bridge (Ross, 1995) is a stochastic process that transits between two fixed endpoints, which is formulated as $X_{t}=W_{t}|(W_{t_{1}},W_{t_{2}})$, where $W_{t}$ is a standard Wiener process with distribution $\mathcal{N}(0,t)$. We can write a Brownian Bridge as $X_{t}=W_{t}|(W_{0},W_{T})$ to define a diffusion process. When $W_{0}=a,W_{T}=b$, we have:
@@ 3.2. Formulation of Diffusion-based VFI
The goal of video frame interpolation is to estimate the intermediate frame $I_{n}$ given the previous frame $I_{0}$ and the next frame $I_{1}$. n is set to 0.5 to interpolate the frame in the middle of $I_{0}$ and $I_{1}$. In latent diffusion models (Rombach et al., 2022), there is an autoencoder that encodes images to latent representations and decodes images from latent representations. The diffusion process denoises a latent representation, and the decoder reconstruct it back to an image. Since the initial noise is random, the decoded images are diverse images when they are sampled repetitively with the same conditions such as poses. Instead of diversity, VFI looks for a deterministic ground truth, which is the intermediate frame. To estimate the ground truth intermediate frame, we only need to estimate the corresponding latent representation in the LDM-based framework. Therefore, LDM-based VFI can be split into two stages: autoencoder and ground truth estimation. The two stages are defined as:
@@
The autoencoder is modeled with VQModel (Rombach et al., 2022) in Section 3.3, and the ground truth estimation is accomplished by our consecutive Brownian Bridge Diffusion in Section 3.4. During inference, both stages are combined as shown in Figure 2 (c), where we decode diffusion-generated latent representation $\hat{\mathbf{x}}$. Via such formulation, we have a novel method to analyze the LDM-based VFI method. If images decoded from $\mathbf{x}$ (Figure 2 (a)) have similar visual quality to images decoded from $\hat{\mathbf{x}}$ (Figure 2 (c)), then the diffusion model achieves a strong performance in ground truth estimation, so it will be good to develop a good autoencoder. On the other way round, the performance of ground truth estimation can be potentially improved by redesigning the diffusion model.